# Submodular Optimization for Keyframe Selection & Usage in SLAM

### v1

@@ VI Results
Our experiments use datasets collected at UCLA with a 10Hz 32-beam Ouster OS0 and at the Army Research Laboratory Graces Quarters facility with a 20Hz 128-beam Ouster OS1 mounted on a Clearpath Warthog. Graces Quarters datasets include a 2.3 km loop and 6.2 km out-and-back shown in Fig. 1 and Fig. 3. The neural network was trained on UCLA campus data using Adam optimizer for 30 epochs.
Online Keyframe Selection. The first experiment compares end-to-end accuracy and memory usage (measured by RSS) of our approach against DLIOM [11]—from which we inherit most of our codebase—and a competitive open-source LIO algorithm [15]. We note that memory allocation is a complex process which RSS does not entirely capture as running these algorithms on different hardware can affect RSS. All results in Table 1 were made using an i7-10875H 8-Core CPU with 16GB of RAM, with test D (Forest Loop) being included to demonstrate the memory demands of long-duration missions. We achieve comparable end-to-end accuracy with DLIOM despite using 64% less memory on average. This is due to a reduction in the number of keyframes selected, as we select 80% less keyframes than DLIOM. Our efficient approach to keyframe selection is shown in the bottom-right of Fig. 1 where eight keyframes each capture unique areas around a large sculpture.
Submodular Submap Generation. The submap generation uses fewer keyframes to improve per scan computation time without affecting scan alignment as measured by the minimum Hessian eigenvalue. Fig. 4 shows a reduction from an average submap size of 10.4 keyframes using DLIOM to 2.3 using our method over a 6.2 km, 43 minute trajectory using an identical set of keyframes (placed every 5m). We achieve a slightly improved alignment constraint over the trajectory, improving from an average minimum Hessian eigenvalue of 377 to 407, while also decreasing the average computation time per scan from 67 to 45 ms. This improvement is due to the submap generation selecting a smaller set of keyframes which can include distant keyframes with nonzero marginal value that heuristics might not consider.

### v2

@@ VI Results
We use a modified version of DLIOM [11] where keyframes used for scan-to-submap alignment are the primary source of memory usage. Our experiments use datasets collected at UCLA with a 10 Hz 32-beam Ouster OS0 and at the Army Research Laboratory Graces Quarters facility with a 20 Hz 128-beam Ouster OS1 mounted on a Clearpath Warthog. Graces Quarters datasets include a 2.3 km loop and 6.2 km out-and-back shown in Fig. 1 and Fig. 3. The neural network was trained on UCLA campus data for 30 epochs.
Online Keyframe Selection. The first experiment compares end-to-end accuracy and memory usage (measured by RSS) of our approach against DLIOM [11] and a competitive open-source LIO algorithm [15]. We note that memory allocation is a complex process which RSS does not entirely capture as running these algorithms on different hardware can affect RSS. All results in Table 1 were made using an i7-10875H 8-Core CPU with 16 GB of RAM, with test D (Forest Loop) being included to demonstrate the memory demands of long-duration missions. We achieve comparable end-to-end accuracy with DLIOM despite using 64% less memory on average. This is due to a reduction in the number of keyframes selected, as we select 80% less keyframes than DLIOM. Our efficient approach to keyframe selection is shown in the bottom-right of Fig. 1 where eight keyframes each capture unique areas around a large sculpture.
Submodular Submap Generation. Submap generation uses fewer keyframes to improve computation time without affecting scan alignment. Fig. 4 shows a reduction from an average submap size of 10.4 keyframes using DLIOM to 2.3 using our method over a 6.2 km, 43 minute trajectory using an identical set of keyframes (placed every 5 m). We achieve a slightly improved alignment constraint over the trajectory, improving from an average minimum Hessian eigenvalue of 377 to 407, while also decreasing the average computation time per scan from 67 to 45 ms. This improvement is due to the submap generation selecting a smaller set of keyframes which can include distant keyframes with nonzero marginal value that heuristics might not consider.